OpenStudy (anonymous):

solve the equation. 2g+14+5g=8g-3+8 g=?????

jimthompson5910 (jim_thompson5910):

Can you simplify 2g+14+5g ?

OpenStudy (anonymous):

add 2g and 5g?

jimthompson5910 (jim_thompson5910):

good to get what?

OpenStudy (anonymous):

7g

jimthompson5910 (jim_thompson5910):

you got it so 2g+14+5g=8g-3+8 becomes 7g+14=8g-3+8

OpenStudy (netlopes1):

2g+14+5g=8g_3+8 -> 2g+5g-8g=-3+8-14 -> -g=-9 -> g=9

jimthompson5910 (jim_thompson5910):

now can you simplify 8g-3+8 ?

OpenStudy (anonymous):

8g-11

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

-3+8 is really 8-3

OpenStudy (anonymous):

8g-5?

jimthompson5910 (jim_thompson5910):

8-3 is positive 5, so it simplifies to 8g+5

jimthompson5910 (jim_thompson5910):

So 7g+14=8g-3+8 then becomes 7g+14=8g+5

jimthompson5910 (jim_thompson5910):

So after simplifying and combining like terms, we get 7g+14=8g+5 What's the next step?

OpenStudy (anonymous):

find what g is? haha

jimthompson5910 (jim_thompson5910):

lol yes that's the ultimate goal, but any ideas on the next step?

OpenStudy (anonymous):

multiply?

jimthompson5910 (jim_thompson5910):

The goal is to get g all by itself, so why don't get all the g terms to one side

jimthompson5910 (jim_thompson5910):

So let's move all the g terms to the right side

OpenStudy (anonymous):

ok so like 14+5=7g+8g ????

jimthompson5910 (jim_thompson5910):

So how do we move that 7g on the left side to the right?

jimthompson5910 (jim_thompson5910):

close...but not quite

OpenStudy (anonymous):

7x8?

jimthompson5910 (jim_thompson5910):

We have 7g+14=8g+5 to move that 7g over, we subtract 7g from both sides Notice how 7g-7g on the left becomes 0g or just 0 (and it goes away) So 7g+14=8g+5 becomes 14=8g+5-7g

jimthompson5910 (jim_thompson5910):

Then we move that "5" over to the left side by subtracting 5 from both sides 14=8g+5-7g 14 - 5=8g+5-7g - 5 14 - 5 = 8g-7g .... Notice how 5 - 5 = 0

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so what do you get after simplifying?

OpenStudy (anonymous):

9=1g?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

or just 9 = g which flips to g = 9

OpenStudy (anonymous):

ok cool thanks!!! again!!! haha

jimthompson5910 (jim_thompson5910):

you're welcome